Factorize:
x^2-x-20=0

Answer by ewatrrr(24785)

 
Hi,
Factorize:
x^2-x-20=0
(x-5)(x+4)= 0
x = 5 0r x = -4
May check factoring Using FOIL
F First terms x^2
O Outside terms 4x
I Inside terms -5x
L Last terms -20